### v3

# Beyond Ensembles: Simulating All-Atom Protein Dynamics in a Learned Latent Space

Abstract
Simulating the long-timescale dynamics of biomolecules is a central challenge in computational science. While enhanced sampling methods can accelerate these simulations, they rely on pre-defined collective variables that are often difficult to identify. A recent generative model, LD-FPG, demonstrated that this problem could be bypassed by learning to sample the static equilibrium ensemble as all-atom deformations from a reference structure, establishing a powerful method for all-atom ensemble generation. However, while this approach successfully captures a system’s probable conformations, it does not model the temporal evolution between them. We introduce the Graph Latent Dynamics Propagator (GLDP), a modular component for simulating dynamics within the learned latent space of LD-FPG. We then compare three classes of propagators: (i) score-guided Langevin dynamics, (ii) Koopman-based linear operators, and (iii) autoregressive neural networks. Within a unified encoder–propagator–decoder framework, we evaluate long-horizon stability, backbone and side-chain ensemble fidelity, and functional free-energy landscapes. Autoregressive neural networks deliver the most robust long rollouts; score-guided Langevin best recovers side-chain thermodynamics when the score is well learned; and Koopman provides an interpretable, lightweight baseline that tends to damp fluctuations. These results clarify the trade-offs among propagators and offer practical guidance for latent-space simulators of all-atom protein dynamics.
1 Signal Processing Laboratory (LTS2), EPFL, Lausanne, Switzerland
2 Institute of Bioengineering, EPFL, Lausanne, Switzerland
3 Ludwig Institute for Cancer Research, Lausanne, Switzerland
Corresponding author: aditya.sengar@epfl.ch
1 Introduction
Molecular simulations are indispensable for studying the complex dynamics that govern biological function, yet brute-force approaches struggle to access the slow, functionally relevant motions—such as protein folding, ligand binding, or allosteric switching—due to rugged energy landscapes and the dominance of rare events [1, 2]. To mitigate this gap—beyond what enhanced sampling can offer when suitable collective variables are hard to specify  [3] —a complementary strategy has gained traction: recasting the simulation challenge from one of brute-force integration to one of representation learning. In this representation-first view, the simulation problem is recast as a modular encoder–propagator–decoder pipeline: an encoder maps high-dimensional atomic configurations into a continuous, low-dimensional latent space; a propagator evolves the system’s state within this simplified space; and a decoder maps the resulting latent trajectory back to all-atom coordinates [4, 5].
Progress in this paradigm follows two complementary directions: The first focuses on learning the underlying physics, employing score-based diffusion  [6, 7], flow matching  [8], and energy-based models  [9] to learn generative surrogates—sometimes yielding differentiable force fields—that implicitly define the system’s potential of mean force [10, 11, 12, 13]. The second centers on learning simplified dynamical coordinates, using time-aware autoencoders or Koopman/DMD analysis  [14, 15] to discover an intrinsic manifold where the long-term dynamics become stable, predictable, or even approximately linear [16, 17, 18, 19, 20].
Despite the promise of latent-space simulation, the choice of propagator—the engine that advances the state—entails a trade-off between physical faithfulness, long-horizon stability, and expressive capacity. We adopt LD-FPG [21] as a fixed encoder–decoder and compare, in the same latent, three propagators: (i) score-guided Langevin dynamics, (ii) a Koopman linear operator, and (iii) an autoregressive neural network (NN). Our analysis reveals a clear trade-off: the MLP is most stable, the score-guided model is most physically detailed for local motions, and the Koopman operator is a simple but overly rigid baseline.
Contributions:
(1) A controlled study that holds the LD-FPG encoder–decoder fixed while swapping the latent propagator, isolating where gains arise.(2) Three propagators within one latent: score-guided Langevin, a Koopman operator, and an autoregressive NN.(3) Benchmarks on alanine dipeptide, 7JFL, and $_{\mbox{1}}A$R.(4) First learned, all-atom rollout (to our knowledge) that recovers a GPCR activation surface ($_{\mbox{2}}A$R).
2 Related Work
The encoder–propagator–decoder blueprint.
Simulating in a low-dimensional space is well established. Molecular Latent Space Simulators (LSS) [4] and Deep Generative MSMs (DeepGenMSM) [5] factorize the task into an encoder that maps coordinates to latent variables, a propagator that advances them, and a decoder that returns all-atom structures. Realizations include autoregressive simulators with recurrent propagators [22], trajectory-level generators that synthesize MD clips in coordinate space [23], LSS variants that target free-energy surfaces and transition kinetics [24], and joint conformation–dynamics models trained via autoregression [25]. We follow this modular design but (i) hold the LD-FPG decoder fixed to control reconstruction quality and (ii) focus our analysis on the choice of latent propagator.
Learning dynamically aware latent spaces (encoders).
The quality of a latent simulation hinges on the quality of the latent space itself. While linear methods like time-lagged independent component analysis (TICA) remain strong baselines for identifying slow variables [26], deep learning has enabled far more expressive encoders. Time-lagged autoencoders such as the Variational Dynamics Encoder (VDE) learn non-linear representations predictive over a time delay $\Delta t$ [16, 27]. VAMPnets use a variational principle to approximate the leading Koopman eigenfunctions, corresponding to the slowest dynamical processes [17]. To remove hand-crafted features, modern approaches leverage graph neural networks (GNNs) to learn CVs directly from coordinates in a permutation- and symmetry-aware manner [28, 20, 29], often with information-bottleneck objectives to explicitly optimize predictiveness of future states [30]. For a recent overview of time-lagged encoders and related autoencoder variants, see [31].
Linear dynamics in latent space (Koopman/DMD propagators).
The Koopman-operator framework provides a route to linearize non-linear dynamics in a learned observable space [18]. Extended dynamic mode decomposition (EDMD) [32] and DMD [15] showed that a wide class of systems admits accurate linear predictors; subsequent work learns such observables end-to-end with deep encoders [14]. This has inspired Koopman autoencoders that enforce a linear evolution rule within the latent space [19, 33, 34], often yielding strong long-horizon stability by reducing compounding errors in nonlinear rollouts. Beyond linear Koopman rules, implicit transfer-operator learning uses a diffusion model to parameterize the Markov kernel at multiple lags, yielding consistent long-horizon rollouts under SE(3) symmetry  [35].
Non-linear sequence models (neural propagators).
A complementary path directly learns the non-linear transition function. The Learning Effective Dynamics (LED) framework uses LSTMs to propagate latent variables and capture memory effects [22]; others pair RNNs with physical resampling such as Maximum Caliber to enforce kinetic consistency [36]. Continuous-time learners (Neural ODEs) offer flexible parametric flows [37], with extensions to stochastic settings via Neural SDEs [38]. Recent work also couples an autoregressive temporal module with an SE(3)-equivariant diffusion decoder to jointly learn conformations and their ordering from MD (ConfRover)  [25]. Our MLP propagator serves as a streamlined baseline in this family.
Physics-guided stochastic models (Langevin and diffusion).
Generative models increasingly embed statistical-mechanics structure. Score-based diffusion has been used to learn effective force fields for coarse-grained MD [10], tying the denoising objective to forces and Langevin-type simulators via the SDE view of score models [6]. For accelerated evolution, frame-to-frame diffusion/score surrogates have been proposed, including DiffMD for trajectory synthesis conditioned on the previous frame [39] and Score Dynamics, which targets picosecond time steps while preserving short-time kinetic signals [12]. Flow-matching has likewise been applied to coarse-grained rollouts without explicit forces (F3low; manifold/SE(3)-equivariant) [13] and to learning coarse-grained free-energy models via density estimation [40]. A recurring theme is consistency between the sampled equilibrium ensemble and the stationary law of the learned dynamics; enforcing Fokker–Planck consistency during training improves long-horizon behavior [11].
GPCRs as a proving ground for slow dynamics.
G protein-coupled receptors (GPCRs) are a demanding testbed for dynamic models, as their function is governed by slow transitions between metastable states [1, 2]. Decades of structural and computational work have characterized the micro-switches and allosteric pathways involved in their activation [41, 42], making them ideal systems for benchmarking. Our evaluation on $_{\mbox{1}}A$R and $_{\mbox{2}}A$R tests the ability of each latent propagator to recover these functionally critical motions.
Positioning among recent generative simulators.
While geometric diffusion models achieve strong results for static 3D structures and poses [43, 44], our focus is the distinct problem of simulating temporal dynamics. Recent efforts include coordinate-space trajectory synthesis with diffusion (GeoTDM) [45], SE(3)-equivariant flow matching for frame-to-frame coarse-grained rollouts (F3low) [13], and neural operators for multi-step 3D dynamics (EGNO) [46]. In parallel, latent and physics-informed approaches aim to accelerate sampling or encode structure—LAST for adaptive MD [47], ConfRover for joint conformation–dynamics learning [25], and NeuralMD, a multi-grained group-symmetric differential-equation model for protein–ligand binding dynamics [48]. Complementing these directions, we isolate the propagator by comparing linear (Koopman), neural, and score-guided dynamics within the same learned latent and fixed decoder, enabling a controlled comparison of long-horizon stability, ensemble fidelity, and functional free-energy surfaces.
3 Methods
Figure 1: Framework overview. A pre-trained LD-FPG encoder (ChebNet; left) maps all-atom coordinates $X(t)$ to a pooled latent $z(t)$. Within the fixed LD-FPG latent, GLDP advances the state via one of three propagators (red box): (a) score-guided Langevin using the LD-FPG denoiser to estimate $s_{\theta}(z,\tau)=\nabla_{z}\log p_{\tau}(z)$ at a fixed low-noise level; (b) an autoregressive NN $z_{t+1}=f_{\theta}(z_{t})$; and (c) a Koopman linear operator $z_{t+1}=Az_{t}$. The frozen LD-FPG decoder (right) maps the latent trajectory back to all-atom structures $\hat{X}(t+\Delta t)$.
GLDP extends the LD-FPG encoder–decoder backbone  [21] (see App. A.2) by adding a temporal propagator that operates entirely within the LD-FPG latent. The encoder–decoder is kept fixed (frozen) in all experiments; only the propagator is swapped.
3.1 GLDP: Propagation in the LD-FPG Latent Space
We adopt the LD-FPG encoder–decoder backbone. The encoder, a Chebyshev Graph Neural Network (ChebNet), maps all-atom coordinates $X(t)\in\mathbb{R}^{N\times 3}$ to a pooled latent $z(t)\in\mathbb{R}^{d}$. Trained on an MD trajectory, the encoder produces a time series $\{z_{0},z_{1},\dots,z_{M}\}$ that captures conformational variation. We then learn a latent-time update of the form
$z_{t+1}\;=\;f(z_{t})\;+\;\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,\sigma_{\eta}^{2}I),$ (1)
where $f$ is the propagator and $\eta_{t}$ is optional rollout noise. Unless noted, training pairs use a one-frame stride $(z_{t},z_{t+1})$, latents are standardized per dimension using training-set statistics before fitting the propagator (and unstandardized before decoding), and the initial state $\hat{z}_{0}$ is the encoding of a held-out MD frame. Key notation appears in Table 10; a glossary is in Table 11.
3.2 Koopman Propagator via Dynamic Mode Decomposition
The Koopman variant approximates latent evolution with a single linear operator:
$z_{t+1}\approx A\,z_{t},\qquad A\in\mathbb{R}^{d\times d}.$
where $A$ is the time-independent Koopman operator that we aim to learn from data. We estimate $A$ with Dynamic Mode Decomposition (DMD) [15]. First, we arrange our latent trajectory into two snapshot matrices, where each column is a state of the system at a point in time,
$\mathbf{X}=[z_{0},\dots,z_{M-2}],\qquad\mathbf{Y}=[z_{1},\dots,z_{M-1}],$
and solve the least-squares map $\min_{A}\|\mathbf{Y}-A\mathbf{X}\|_{F}^{2}$, yielding
$A\;=\;\mathbf{Y}\,\mathbf{X}^{+},$
where $\mathbf{X}^{+}$ is the Moore–Penrose pseudoinverse. For numerical stability and to regularize the model, the pseudoinverse is computed using Singular Value Decomposition (SVD). By truncating the SVD to a lower rank $r$, we can filter out noise and focus the model on the most energetically significant dynamic modes. Once learned, the operator $A$ is used to generate new trajectories autoregressively from an initial state $\hat{z}_{0}$; a small Gaussian rollout noise $\eta_{t}\sim\mathcal{N}(0,\sigma_{\text{koop}}^{2}I)$ is added to match the one-step residual variance. Implementation details are provided in App. A.3.
3.3 Autoregressive Neural Network Propagator
To learn the latent transition directly, we use an autoregressive NN. The model parameterizes a deterministic nonlinear map
$z_{t+1}=f_{\theta}(z_{t}),$ (2)
trained with a one–step objective over pairs $(z_{t},z_{t+1})$:
$\mathcal{L}(\theta)\;=\;\frac{1}{M-1}\sum_{t=0}^{M-2}\big\|f_{\theta}(z_{t})-z_{t+1}\big\|^{2}.$ (3)
To reduce error accumulation in long rollouts, we use dropout and weight decay (see App. A.4 for architecture and training settings). After training, trajectories are generated by iterative application from an initial state $\hat{z}_{0}$:
$\hat{z}_{i+1}\;=\;f_{\theta}(\hat{z}_{i})\;+\;\eta_{i},\qquad\eta_{i}\sim\mathcal{N}(0,\sigma_{\text{roll}}^{2}I),$ (4)
where $\sigma_{\text{roll}}$ is chosen on a validation split to match the one–step latent residual variance of $z_{t+1}-z_{t}$. (Unless noted, $\hat{z}_{0}$ is the encoding of a held-out MD frame and latents are standardized per dimension before fitting and unstandardized before decoding.) The network architecture, training procedure, and rollout settings are detailed in Appendix A.4.
3.4 Score-Guided Langevin Propagator
This propagator uses a learned score of the latent equilibrium to define a stochastic simulation. The procedure has two stages.
Stage 1: score model.
We train a time-conditional denoiser $\epsilon_{\theta}(z_{t},t)$ on latent embeddings with a variance-preserving (VP) schedule as in DDPMs. The forward corruption is
$z_{t}\;=\;\sqrt{\bar{\alpha}_{t}}\,z_{0}\;+\;\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),\quad\bar{\alpha}_{t}\!=\!\prod_{s=1}^{t}(1-\beta_{s}).$
When the network is trained to predict noise, the perturbed-data score is
$s(z_{t})\;=\;\nabla_{z}\log p_{t}(z_{t})\;\approx\;-\frac{\epsilon_{\theta}(z_{t},t)}{\sqrt{1-\bar{\alpha}_{t}}}\,.$
To approximate the unperturbed score $s(z)$ at test time, we query a low noise level $t_{\text{noise}}\!\approx\!0$ (we use $t_{\text{noise}}{=}0$ unless stated).
Stage 2: Langevin simulation in latent space.
We integrate an overdamped Langevin SDE with Euler–Maruyama, using the score as the drift:
$z_{t+1}\;=\;z_{t}\;+\;T\,\Delta t\,s(z_{t})\;+\;\sqrt{2\,T\,\Delta t}\;\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,\sigma_{\eta}^{2}I).$ (5)
The effective diffusion variance per step is $2\,T\,\Delta t\,\sigma_{\eta}^{2}$; we set $\sigma_{\eta}{=}1$ unless otherwise noted. Equivalently, if $p(z)\!\propto\!\exp(-U(z)/T)$, then $T\,s(z)=-\nabla U(z)$, so the drift corresponds to the gradient of an effective energy surface in latent space.
The temperature $T$, step size $\Delta t$, and other implementation details, such as score clipping and sampling stride are further described in Appendix A.5.
Target ensemble.
Our propagators evolve only the latent coordinate $z$ (no momenta, no barostat/box dynamics). The objective is to reproduce the configuration-space equilibrium of the reference MD at its thermodynamic state. Let $p_{X}(x)$ denote the equilibrium distribution of all-atom coordinates $x$ from the source MD, and let $E[\cdot]$ and $D[\cdot]$ be the fixed LD-FPG encoder and decoder. The target latent law is the pushforward $p_{Z}\;=\;E_{\#}\,p_{X}\quad\text{with}\quad z=E(x)$ and decoded samples are $\hat{x}=D(z)$. We fit propagators so that the stationary decoded trajectory $\{\hat{x}(t)\}$ matches reference statistics in $X$-space (e.g., dihedral histograms, RMSF, contacts, and TM-distance free-energy surfaces). Because no Hamiltonian with momenta is integrated in latent space, we do not claim NVE energy conservation or exact physical kinetics. The score-guided Langevin propagator targets a canonical law in latent space whose invariant density is $p_{Z}$ when the score is exact; Koopman and the autoregressive NN model approximate the one-step transition kernel $K(z_{t+1}\!\mid z_{t})$ observed from $(z_{t},z_{t+1})$ under the fixed encoder. Unless stated otherwise, comparisons of timescales are diagnostic rather than calibrated to wall-clock time.
4 Results and Discussion
Key result. GLDP operates reliably across scales (ADP $\rightarrow$ 7JFL $\rightarrow$ $_{\mbox{1}}A$R) and reproduces the $_{\mbox{2}}A$Ractivation surface; this cross-scale pattern is the main message of the paper. See Appendix A.7 for $_{\mbox{1}}A$R ablations (encoder width, pooling size, and Langevin/denoiser settings).
We first position GLDP against recent generative simulators using GLDP-NN (GLDP with the autoregressive NN propagator)(Sec. 4.1). We then compare Koopman, NN, and Langevin within the same latent space on ADP and  $_{\mbox{1}}A$Rfor stability and ensemble fidelity (Sec. 4.2, 4.3), and finally assess all three on  $_{\mbox{2}}A$Rfunctional surfaces (Sec. 4.4).
4.1 Comparative Benchmarks on Trajectory Generation
To benchmark GLDP, we compared it against a diverse set of state-of-the-art models: the conceptually similar Latent Space Simulator (LSS) [4], the video-synthesis-based MD-Gen [23], and GeoTDM [45], a leading coordinate-based diffusion model. All methods were evaluated on equilibrium fidelity, flexibility, and kinetics across systems of increasing complexity: the canonical alanine dipeptide (ADP), the 7JFL protein, and the functionally complex $_{\mbox{1}}A$R GPCR. The quantitative results are summarized in Table 1.  [^1] GeoTDM on 7JFL and $_{\mbox{1}}A$R exceeded 80 GB GPU memory (H100). MD-Gen does not support capped termini in our ADP setup; ADP results omitted.
Table 1: Benchmark of dynamics models. GLDP is compared against LSS, MD-Gen, and GeoTDM. Lower JSD is better. $B$: backbone; SC: sidechain. For RMSF, predicted values closer to the ground-truth row are better. Higher contact correlation is better. Bold indicates the best result among models. N/A: not applicable.
JSD Avg. RMSF (Å)
System Model Dihedral (BB/SC) Coord. (BB/SC) Contact Corr. Backbone Sidechain
ADP Ground Truth N/A N/A N/A 0.56 1.34
LSS 0.016 / N/A 0.621 / 0.633 N/A 0.54 0.90
GeoTDM 0.123 / N/A 0.444 / 0.395 N/A 0.43 0.77
GLDP 0.061 / N/A 0.366 / 0.386 N/A 0.61 0.89
7JFL Ground Truth N/A N/A N/A 2.09 3.05
LSS 0.027 / 0.097 0.691 / 0.682 0.935 0.60 0.78
MD-Gen 0.035 / 0.090 0.377 / 0.361 0.960 0.99 1.73
GLDP 0.037 / 0.081 0.426 / 0.459 0.949 1.44 2.30
$_{\mbox{1}}A$R Ground Truth N/A N/A N/A 1.49 1.93
LSS 0.134 / 0.146 0.657 / 0.602 0.971 1.11 1.41
MD-Gen 0.033 / 0.088 0.106 / 0.117 0.941 0.86 1.31
GLDP 0.019 / 0.067 0.063 / 0.087 0.985 1.22 1.48
Ensemble Fidelity:
We assess equilibrium fidelity via Jensen–Shannon divergence (JSD) over global coordinates and local dihedrals. On the small ADP system, GLDP attains the best coordinate JSD, while LSS edges out on backbone dihedral JSD and is closest to ground-truth RMSF. On the medium 7JFL system, MD-Gen leads on coordinate JSD and contact correlation, LSS is best on backbone-dihedral JSD, and GLDP is best on side-chain–dihedral JSD and most closely matches backbone/side-chain RMSF amplitudes. On the large GPCR $_{\mbox{1}}A$R, GLDP achieves the lowest JSDs for both coordinates and dihedrals, the highest contact correlation, and the closest RMSFs, indicating the most faithful recovery of both tertiary structure and local conformational statistics. Overall, GLDP provides the strongest balance across metrics on complex proteins, while MD-Gen and LSS retain advantages on specific axes for smaller or mid-scale systems.
Molecular flexibility (RMSF).
Matching the amplitude of motion is as important as matching mean structure, so we compare backbone/side-chain RMSF to the reference MD. On ADP, all methods undershoot side-chain flexibility; GLDP slightly overestimates backbone RMSF (0.61 Å vs. 0.56 Å), while LSS is the closest overall on ADP (0.54/0.90 Å). On 7JFL, GLDP best recovers fluctuation magnitudes for both backbone and side chains (1.44/2.30 Å vs. 2.09/3.05 Å), whereas MD-Gen (0.99/1.73 Å) and LSS (0.60/0.78 Å) are overly rigid. On $_{\mbox{1}}A$R, all models are conservative, but GLDP remains closest to the reference variability (1.22/1.48 Å vs. 1.49/1.93 Å). In sum, GLDP most faithfully reproduces fluctuation amplitudes on the larger proteins, complementing its strong ensemble statistics.
Global contacts / long-horizon stability:
As a coarse proxy for long-timescale stability, we compare the Pearson correlation between model and reference time-averaged contact maps (Appendix A.6). Contact patterns are well preserved by the top models (Pearson $r>0.93$ where evaluated). On 7JFL, MD-Gen best matches the contacts ($r=0.960$), while GLDP attains the highest correlation on $_{\mbox{1}}A$R ($r=0.985$) (ADP omitted: system too small for a meaningful contact map.)
Summary.
Across flexibility and kinetics, GLDP offers an excellent balance of realism and stability. On the complex $_{\mbox{1}}A$R receptor, it is the top performer across all metrics, achieving the lowest Jensen-Shannon Divergence for both coordinates and dihedrals, the highest contact correlation ($r=0.985$), and the most accurate fluctuation amplitudes (RMSF). This strong performance extends to other systems, where GLDP consistently places as either the best-performing model—such as for coordinate fidelity on ADP and side-chain distributions on 7JFL—or as a close competitor to the leader. Crucially, GLDP faithfully matches the flexibility of larger proteins where competing methods are often overly rigid, positioning it as a highly dependable latent-space simulator.
4.2 Long-horizon stability: autoregressive NN is most robust
Figure 2: Stability over long rollouts. RMSD and lDDT versus frame index for (a) alanine dipeptide and (b) $_{\mbox{1}}A$R. We define failure time as the first frame whose lDDT (relative to the initial frame) drops below 0.65. On $_{\mbox{1}}A$R, the autoregressive NN remains stable for the entire 10,000-frame horizon (no failure), while Langevin and Koopman fail earlier; on alanine dipeptide, Koopman and NN persist for thousands of frames whereas Langevin fails early.
Figure 2 tracks RMSD/lDDT [^2] The local Distance Difference Test (lDDT) evaluates local structural quality by measuring the preservation of inter-atomic distances. over rollout length and reports the failure time (first frame with lDDT <0.65<0.65 relative to the initial frame). On alanine dipeptide, Koopman and the autoregressive NN remain stable for thousands of frames (failure at 4,443 and 3,176, respectively), whereas Langevin terminates early (206). On $_{\mbox{1}}A$R, the NN completes the full 10,000-frame horizon without failure; Langevin reaches 7,476 and Koopman 5,740. Taken together, the NN is the most robust propagator for long-horizon rollouts on both a small peptide and a large GPCR, with Koopman providing a competitive linear baseline and Langevin exhibiting sensitivity to score/integration settings.
What a “frame” means for Langevin (and how we calibrated it).
For Koopman and NN, one frame equals one dataset stride ($z_{t}\!\to\!z_{t+1}$), so the horizontal axis matches the MD sampling stride. The Langevin propagator integrates a latent SDE with internal step size $\Delta t$ and a separate sampling stride $s$; a plotted “frame” is a sampled SDE state, not a single integrator step. To make curves comparable, we calibrated $(\Delta t,s)$ to match the short-horizon RMSD-per-frame of the reference MD (Sec. A.5): for alanine, $\Delta t=10^{-8}$ and s=1s{=}1; for $_{\mbox{1}}A$R, $\Delta t=10^{-10}$ and s=20s{=}20 (effective time per plotted frame $s\,\Delta t$). Reported Langevin failure indices therefore count sampled outputs.
Why Langevin fails earlier on alanine.
We attribute the early failure on alanine to three factors: (i) a sparsely sampled latent manifold from a large MD frame stride, yielding a rough score estimate $s(z)=\nabla_{z}\log p(z)$; (ii) larger apparent per-frame displacements in a tiny system (RMSD $\sim$0.9 Å) making the lDDT threshold stringent; and (iii) step sizes tuned for GPCRs that are aggressive for alanine under an imperfect score, pushing the SDE off-manifold. In $_{\mbox{1}}A$R, motions distribute over many atoms (lDDT penalizes less), and subsampling (s=20s{=}20) further smooths per-frame jumps.
4.3 Fidelity to the Equilibrium Ensemble
Table 2: Dihedral free-energy fidelity (JSD; lower is better). Values computed on normalized 2D densities with periodic binning.
[…]
Koopman 0.138 0.043 0.121
Autoreg. NN 0.061 0.019 0.067
Langevin 0.164 0.052 0.058
We quantify ensemble fidelity by comparing free–energy surfaces (FES) of backbone $(\phi,\psi)$ and side–chain $(\chi_{1},\chi_{2})$ dihedrals (Fig. 3). The Jensen–Shannon divergence (JSD) between the normalized model and reference densities serves as our primary error metric (lower is better), with all results summarized in Table 2.
Backbone $(\phi,\psi)$ fidelity is dominated by the autoregressive NN.
On backbone dihedrals, the autoregressive NN consistently outperforms the other propagators. For alanine dipeptide, the NN achieves the lowest divergence (JSD =0.061=0.061), accurately capturing the canonical Ramachandran basins; the Koopman model’s linear rigidity yields a markedly less accurate landscape (JSD =0.138=0.138). This trend is even more pronounced on $_{\mbox{1}}A$R, where the NN again produces the most faithful backbone surface (JSD =0.019=0.019), while both Koopman and Langevin lag behind.
Side-chain $(\chi_{1},\chi_{2})$ rotamers favor the Langevin propagator.
The picture reverses for side-chain rotamers on $_{\mbox{1}}A$R. Here, the score-guided Langevin propagator is superior, yielding the sharpest rotamer bands and the lowest JSD ($0.058$). This likely reflects its thermodynamically consistent drift (Ts(z)T\,s(z)) plus isotropic noise, which excels at sampling sharp, multi-modal distributions corresponding to distinct rotameric states. The NN remains close (JSD =0.067=0.067) but slightly oversmooths high-barrier regions, whereas Koopman merges them.
Taken together, these results highlight a trade-off: the autoregressive NN excels at capturing broad backbone dynamics, while the score-guided Langevin propagator is superior for resolving fine-grained, multi-modal side-chain states.
Figure 3: Ensemble fidelity in dihedral space. Free-energy maps for backbone $(\phi,\psi)$ and, for $_{\mbox{1}}A$R, side-chain $(\chi_{1},\chi_{2})$. Alanine dipeptide (left) shows canonical basins recovered by all methods; the Autoregressive NN aligns best with the reference basin shapes. For $_{\mbox{1}}A$R(middle/right), the NN most closely matches backbone structure, while the score-guided Langevin propagator recovers the sharpest rotamer bands for side-chains.
4.4 GPCR function: activation surface in $_{\mbox{2}}A$R
Figure 4: Functional free–energy surface for $_{\mbox{2}}A$R. Left: 2D free energy over TM3–6 and TM3–7 distances (reference heatmap; model contours overlaid). Right: 1D profiles along each coordinate. Langevin spans the diagonal valley and its shoulders; the autoregressive NN follows the same valley with a tighter footprint; Koopman centers the basin but amplifies apparent barriers and reduces anisotropy. Free-energy surfaces for each propagator are averaged over 5 independent replicates.
[…]
$_{\mbox{2}}A$R: two-dimensional surface.
The reference surface forms a diagonal valley, indicating coordinated changes in TM3–6 and TM3–7 during activation. Langevin covers this valley most extensively, including the transition corridor connecting inactive and active-like regions. The Autoregressive NN follows the same valley with a tighter footprint, giving good alignment near the basin center but under-sampling the flanks. Koopman identifies the central basin yet shows stiffer, more isotropic contours; barriers appear sharper and the diagonal anisotropy is reduced, which limits coverage along the transition path.
One-dimensional slices.
Along TM3–7 (top-right), Langevin and the NN track the position of the principal minimum and its curvature, with Langevin retaining broader support across the valley shoulders. Koopman exhibits a narrow well and steeper rise away from the minimum, consistent with its tendency to compress variance on the 2D map. Along TM3–6 (bottom-right), Langevin again follows the reference profile across both flanks, while the NN stays close near the minimum but with a slightly skewed tail. Koopman produces a compact basin and elevated shoulders, which reduces sampling of the transition corridor toward larger TM3–6 distances limiting its exploration of the activation corridor and showing that the non-linear propagators are more effective at tracing the complete activation pathway.
4.5 Outlook
Our comparison reveals a fundamental trade-off between the stability of data-driven models and the fine-grained physical accuracy of score-guided methods. This points toward several promising future directions. First is the development of hybrid propagators that could leverage neural networks for stable, long-horizon integration of slow collective motions while using a score-guided Langevin update for local, high-frequency side-chain sampling to ensure thermodynamic correctness. Second, a key avenue for improvement is the end-to-end training of the encoder and propagator to discover a latent representation optimized for stable temporal evolution. Such a latent space could be made more robust by enforcing thermodynamic consistency constraints or by explicitly training the model with a denoising objective, where it learns to predict the original protein structure from an input dataset to which noise has been added. For greater realism, models require rigorous kinetic calibration to match simulation or experimental timescales and should incorporate stronger physical priors like symmetry-aware (SE(3)) architectures, though their computational cost remains a significant challenge for large proteins.
Acknowledgments
A.S. was supported by a grant from the Center of Intelligence Systems (EPFL) to P.B. and P.V. This work was also supported by the Institute of Bioengineering (EPFL), the Signal Processing Laboratory (LTS2, EPFL), the Ludwig Institute for Cancer Research, and the Swiss National Science Foundation (grants 31003A_182263 and 310030_208179).
Appendix A Appendix
This Supplementary Information provides detailed descriptions of the models, training procedures, and evaluation metrics used in this work to ensure reproducibility.
A.1 Simulation Datasets and Code availability
The complete code, containing the pipeline to train the LD-FPG model and run GLDP, is publicly available at: https://github.com/adityasengar/GLDP. The latent trajectories used to train the propagators were derived from three publicly available Molecular Dynamics (MD) simulation datasets. For each system, the original source provided a structure file (PDB) and a coordinate trajectory (XTC). These files were then processed using the LD-FPG framework’s preprocessing scripts to generate the inputs for GLDP, including a condensed.json file that provides consistent, zero-based atom indexing and defines the atom quadruplets required for calculating all backbone and side-chain dihedral angles.
Alanine Dipeptide.
This dataset features N-acetyl-L-alanine-N’-methylamide, a 22-atom molecule commonly known as alanine dipeptide. It is a canonical benchmark for developing and testing new simulation methods due to its simple yet non-trivial conformational landscape, which is primarily described by its two backbone dihedral angles ($\phi,\psi$). The data was sourced from the CMB data repository at ftp.imp.fu-berlin.de and consists of a 250 ns simulation trajectory with solvent molecules removed.
7JFL (ATLAS).
We use entry 7jfl_C from the ATLAS database: dsimb.inserm.fr/ATLAS/…/7jfl_C. ATLAS provides standardized all-atom, explicit-solvent MD trajectories across proteins under a common protocol.
[…]
Human $_{\mbox{1}}A$Rtrajectories were taken from the study  [49]. These are explicit-solvent, membrane-embedded GPCR simulations as described in the paper’s Methods/SI and associated data records. Before encoding, we remove solvent and lipids and operate at fixed box; we assess configuration-space statistics after decoding. Where specific thermostat/barostat and T/P values are required, we defer to the original Methods/SI.
[…]
$_{\mbox{2}}A$Rtrajectories come from the  [49]. As above, simulations are explicit solvent with a membrane environment; solvent and lipids are stripped before encoding. We match configuration-space statistics to the reference ensemble and refer to the paper’s Methods/SI for the exact simulation state (e.g., NVT/NPT, temperature, pressure, ions, water model, force field).
A.2 LD-FPG based Encoder-Decoder Backbone
GLDP follows an encoder-propagator-decoder blueprint. We use a fixed, pre-trained encoder-decoder from the LD-FPG model [21] and focus on comparing propagators within its learned latent space.
The latent space for this study was generated using the original LD-FPG framework. The core components are:
  • Encoder: A 4-layer Chebyshev Graph Neural Network (‘HNO‘ model) that maps all-atom coordinates $X(t)\in\mathbb{R}^{N\times 3}$ to per-atom embeddings $Z(t)\in\mathbb{R}^{N\times d_{z}}$.
  • Pooling and Decoder: A ‘ProteinStateReconstructor2D‘ model performs pooling on the per-atom embeddings to generate a single latent vector $z(t)\in\mathbb{R}^{d_{\text{flat}}}$ for the entire conformation. The decoder portion of this model maps this latent vector back to all-atom coordinates.
The script ‘chebet_blind.py‘ was used to train this autoencoder and export the time-series of latent vectors $\{z_{0},z_{1},\dots,z_{M}\}$ into an HDF5 file (‘latent_reps/pooled_embedding.h5‘), which serves as the ground-truth trajectory for training the propagators. Key hyperparameters for the backbone are detailed in Table 3.
Table 3: Configuration of the Pre-trained LD-FPG Backbone Model.
Component Parameter Value
Encoder (HNO)
   Hidden Dimension ($d_{z}$) 16
   Chebyshev Order ($K$) 4
   Graph Construction $k$-Nearest Neighbors ($k=4$)
Decoder (Blind Pooling)
   Pooling Type Blind
   Pooling Output Size ($H\times W$) $50\times 2$
   Flattened Latent Dimension ($d_{\text{flat}}$) 100  (from $50\times 2$)
   MLP Hidden Layers 12
   MLP Hidden Dimension 128
Encoder (HNO).
HNO denotes a Chebyshev graph neural operator: a ChebNet with polynomial filters of order $K$ on the molecular graph (built by kkNN).
Pooling.
“Blind pooling” maps per–atom embeddings $Z(t)\in\mathbb{R}^{N\times d_{z}}$ to a fixed $H\times W$ latent by a learned MLP that aggregates across atoms without explicit residue topology; the decoder inverts this map to all-atom coordinates.
A.3 Koopman Operator Implementation Details
The Koopman operator $A$ was computed using the fit_dmd function, which implements the standard Dynamic Mode Decomposition (DMD) algorithm. The procedure is summarized in Algorithm 1. This method provides a numerically robust way to find the best-fit linear operator $A$ by leveraging the Singular Value Decomposition (SVD) to compute the pseudoinverse of the snapshot matrix $\mathbf{X}$.
Algorithm 1 Computing the Koopman operator with (truncated) DMD
0: Latent snapshots $\{z_{t}\}_{t=0}^{M-1}$, rank cap $r$ (optional), tolerance $\varepsilon$ (optional).
1: Construct $X=[z_{0},\dots,z_{M-2}]$, $Y=[z_{1},\dots,z_{M-1}]$.
2: Compute SVD: $X=U\Sigma V^{\top}$.
3: (Optional) choose rank $r$ by energy or cap; set $U_{r},\Sigma_{r},V_{r}$ to the top-$r$ factors.
4: Form $\Sigma_{r}^{+}$ by inverting only singular values $\varepsilon$ (set others to zero).
5: Compute $X^{+}\leftarrow V_{r}\Sigma_{r}^{+}U_{r}^{\top}$.
6: Set $A\leftarrow YX^{+}$.
6: Linear propagator AA (use $z_{t+1}=Az_{t}$, optionally with small Gaussian rollout noise).
In our implementation, the SVD truncation rank $r$ can be set as a hyperparameter to de-noise the dynamics. If no rank is specified, the full rank of the matrix is used.
A.3.1 Koopman Propagator Hyperparameters
The SVD truncation rank ($r$) and the standard deviation of the Gaussian rollout noise ($\sigma_{\text{koop}}$) are key hyperparameters. They were selected independently for each system to ensure stable, long-horizon rollouts, with the final values reported in Table 4.
Table 4: Koopman model hyperparameters used for each molecular system.
System SVD Truncation Rank ($r$) Rollout Noise Std. ($\sigma_{\text{koop}}$)
$_{\mbox{1}}A$R 10 $2.00\times 10^{-4}$
Alanine Dipeptide 10 $2.60\times 10^{-3}$
A.4 Autoregressive Neural Network Details
The neural network propagator $f_{\theta}$ is implemented as a Multi-Layer Perceptron using PyTorch, as detailed in the neural_propagator.py script. The training and rollout procedures are outlined in Algorithm 2, and the specific hyperparameters are summarized in Table 5. The model is trained on a fixed percentage of the latent trajectory, with the remainder used for validation to select the best-performing model checkpoint.
Table 5: Hyperparameters for the Autoregressive NN Propagator.
Parameter Value / Description
Architecture
   Model Type MLP (PropagatorMLP)
   Input Dimension ($d_{\text{flat}}$) 100
   Hidden Dimension 512
   Hidden Layers 2
   Activation Function ReLU
   Dropout Rate 0.1
Training
   Loss Function Mean Squared Error (MSE)
   Optimizer Adam
   Learning Rate $1\times 10^{-4}$
   Batch Size 128
   Epochs 400
   Frame Skip ($n$ in $z_{t+n}$) 1
   Training Fraction 0.8 (80% train, 20% validation)
   Weight Decay $1\times 10^{-4}$
Algorithm 2 Autoregressive NN Training and Rollout
0: Latent trajectory $\{z_{0},\dots,z_{M}\}$, model $f_{\theta}$, learning rate $\alpha$.
1:// Training
2: Create dataset of pairs $\mathcal{D}=\{(z_{t},z_{t+1})\}_{t=0}^{M-2}$.
3: Initialize model parameters $\theta$.
4:for each epoch do
5:  for each batch $(z_{t},z_{t+1})$ in $\mathcal{D}$ do
6:   Compute loss: $\mathcal{L}\leftarrow\|z_{t+1}-f_{\theta}(z_{t})\|^{2}$.
7:   Update parameters via backpropagation: $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}$.
8:  end for
9:end for
10:
11:// Rollout
12: Initialize trajectory: $\hat{z}_{0}\leftarrow z_{0}$.
13:for $t=0,\dots,M-2$ do
14:  $\hat{z}_{t+1}=f_{\theta}(\hat{z}_{t})+\eta_{t},\quad\eta_{t}\sim\mathcal{N}(0,\sigma_{\text{roll}}^{2}I)$.
15:end for
15: Trained model $f_{\theta}$, generated trajectory $\{\hat{z}_{t}\}$.
We choose $\sigma_{\text{roll}}$ on a validation split to match the one-step latent variance of $z_{t+1}-z_{t}$.
The standard deviation of the rollout noise ($\sigma_{\text{roll}}$) was calibrated for each system by matching the one-step latent residual variance on a validation set, as described in the main text. The specific values used for each simulation are listed in Table 6.
Table 6: System-specific rollout noise for the Autoregressive NN propagator.
System Rollout Noise Std. ($\sigma_{\text{roll}}$)
$_{\mbox{1}}A$R $1.00\times 10^{-3}$
Alanine Dipeptide $1.00\times 10^{-2}$
7JFL $1.50\times 10^{-3}$
A.5 Score-Guided Langevin Propagator Details
The implementation of the Score-Guided Langevin Propagator consists of two main stages.
A.5.1 Denoiser Model Training
First, a time-conditional denoising model, $\epsilon_{\theta}(z_{\tau},\tau)$, is trained on the static set of latent embeddings from the MD trajectory. This model learns to predict the noise that was added to a "clean" sample $z_{0}$ to produce a noisy sample $z_{\tau}$ at diffusion step $\tau$. The training objective is to minimize the Mean Squared Error between the predicted noise and the true noise: $\mathcal{L}=\mathbb{E}_{t,z_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{\tau}}z_{0}+\sqrt{1-\bar{\alpha}_{\tau}}\epsilon,\tau)\|^{2}$. The key hyperparameters for this training stage are listed in Table 7.
Table 7: Hyperparameters for Denoiser Model Training.
Parameter Value / Description
Denoiser Model
   Model Type Time-conditional MLP (DiffusionMLP_v2)
   Input Dimension ($d_{\text{flat}}$) 100
   Hidden Dimension 1024
Diffusion Schedule
   Diffusion Steps ($T_{\text{diff}}$) Range explored: 450 – 1600
   Beta Schedule Linear
   $\beta_{\text{start}}$ Range explored: $5\times 10^{-6}$$0.005$
   $\beta_{\text{end}}$ Range explored: 0.020.020.110.11
Training
   Optimizer Adam
   Learning Rate $1\times 10^{-5}$
   Batch Size 64
   Epochs 50,000
A.5.2 Langevin Simulation
We approximate the score of the unperturbed data distribution by evaluating the model at a low, fixed noise level, $\tau_{\text{noise}}$ (e.g., $\tau=0$). The simulation then proceeds according to Algorithm 3, with simulation-specific hyperparameters listed in Table 8.
Table 8: Hyperparameters for Langevin Simulation.
Parameter Value / Description
Integration Time Step ($\Delta t$) Typical range: $10^{-11}$$10^{-8}$ (arbitrary units)
Temperature ($T$) Typical range: 0.001 – 1.0 (arbitrary units)
Noise Level Timestep ($\tau_{\text{noise}}$) 0 (Evaluates score at lowest noise)
Score Clipping Norm Optional max norm to prevent instability
Algorithm 3 Score-Guided Langevin Simulation
0: Pre-trained denoiser $\epsilon_{\theta}$, initial state $z_{0}$, time step $\Delta t$, temperature $T$, simulation steps $N_{steps}$.
1: Fix a low noise level for score evaluation: $\tau\leftarrow\tau_{\text{noise}}$.
2: Look up noise standard deviation for $\tau$: $\sigma_{\tau}$.
3: Initialize trajectory: $\mathcal{Z}\leftarrow\{z_{0}\}$.
4:for $t=0,\dots,N_{steps}-1$ do
5:  // 1. Compute Score
6:  Predict noise for the current state $z_{t}$: $\epsilon_{pred}\leftarrow\epsilon_{\theta}(z_{t},\tau)$.
7:  Calculate the score: $s(z_{t})\leftarrow-\epsilon_{pred}/\sigma_{\tau}$.
8:  (Optional) Clip score: if $\|s(z_{t})\|>s_{max}$, then $s(z_{t})\leftarrow s(z_{t})\cdot s_{max}/\|s(z_{t})\|$.
9:
10:  // 2. Update State
11:  Calculate drift: $\Delta z_{\text{drift}}\leftarrow T\Delta t\cdot s(z_{t})$.
12:  Calculate diffusion: $\Delta z_{\text{diff}}\leftarrow\sqrt{2T\Delta t}\cdot\mathcal{N}(0,I)$.
13:  Update state: $z_{t+1}\leftarrow z_{t}+\Delta z_{\text{drift}}+\Delta z_{\text{diff}}$.
14:  Append $z_{t+1}$ to $\mathcal{Z}$.
15:end for
15: Generated trajectory $\mathcal{Z}$.
As noted in the main text, key parameters such as the temperature ($T$), time step ($\Delta t$), and sampling stride ($s$) were calibrated for each system to ensure physically realistic and stable trajectories. The specific values used for each protein simulation are provided in Table 9.
Table 9: System-specific simulation parameters for the Score-Guided Langevin propagator. Temperature ($T$) and time step ($\Delta t$) are in latent units.
System Temperature ($T$) Time Step ($\Delta t$) Total SDE Steps Sampling Stride ($s$)
Alanine Dipeptide 5.0 $1.0\times 10^{-8}$ 10,000 1
$_{\mbox{1}}A$R 0.8 $1.0\times 10^{-10}$ 100,000 20
$_{\mbox{2}}A$R 0.8 $1.0\times 10^{-10}$ 100,000 20
Implementation and Calibration Details.
At each step of the Langevin simulation, we evaluate the score $s(z_{t})$ by calling the denoiser $\epsilon_{\theta}(z_{t},t_{\text{noise}})$ and dividing by $\sqrt{1-\bar{\alpha}_{t_{\text{noise}}}}$. To prevent numerical instability from rare large updates, we clip the score’s L2-norm to a maximum value of $c=10$. We optionally subsample the simulated path by emitting one decoded frame every $s$ SDE steps (sampling stride) to control the apparent per-frame displacement. The temperature $T$ and step size $\Delta t$ are critical hyperparameters, which we calibrate for each system to match the short-horizon decoded RMSD-per-frame of the reference MD trajectory (as described in Section 4.2). Additional training and rollout settings, including the diffusion schedule and optimizer, are also detailed in this section.
A.6 Contact map Pearson correlation
[…]
$C_{ij}\;=\;\mathbf{1}\!\left[\bar{D}_{ij}<\delta\right],\qquad C_{ii}=0\;\;\forall i.$
We construct $C^{\mathrm{ref}}$ and $C^{\mathrm{mod}}$ with the same cutoff and residue selection. The reported score is the Pearson correlation coefficient between the flattened matrices:
$r\;=\;\mathrm{corr}\!\left(\mathrm{vec}\!\left(C^{\mathrm{ref}}\right),\mathrm{vec}\!\left(C^{\mathrm{mod}}\right)\right),$
Implementation notes. The procedure uses MDAnalysis for I/O and atom selection (name CA); distances are computed via pdist and squareform, then averaged over frames prior to thresholding. No exclusion of near-sequence pairs (|ij|<3|i-j|<3) is applied in this metric.
A.7 Ablation Studies on $_{\mbox{1}}A$R
We probe three levers that materially affect $_{\mbox{1}}A$R: (i) encoder per-atom width $d_{z}$, (ii) pooled latent size prior to decoding (with fixed W=2W{=}2 so $d_{\text{flat}}{=}2H$), and (iii) Langevin/denoiser settings. We report (i) training-time losses at the level where the lever acts, and (ii) downstream accuracy via backbone/side-chain $\Delta\!\mathrm{RMSF}$ (mean absolute error, Å) and lDDT failure time (first frame with lDDT<0.65<0.65; larger is better). For long-rollout statistics we keep the autoregressive NN and Koopman unchanged and vary only the component under study.
Conventions.
For Langevin rollouts, the sampling stride is s=1s{=}1 (one SDE step per frame). We choose $\Delta t$ by matching the average per-frame RMSD of the decoded trajectory to the MD reference over K=100K{=}100 frames (Sec. 4.2). Unless stated: denoiser query $\tau{=}0$, drift clipping $\|s(z)\|\leq c{=}10$.
A.7.1 (i) Encoder per-atom width $d_{z}\in\{8,16,32\}$
Increasing $d_{z}$ lowers the autoencoder validation MSE as expected, and modestly improves downstream flexibility metrics, with diminishing returns beyond $d_{z}{=}16$.
[…]
8 0.020 0.34 0.54 NN >10,000>\!10{,}000  |  Lang 6,800  |  Koop 5,200
16 0.006 0.26 0.39 NN >10,000>\!10{,}000  |  Lang 7,400  |  Koop 5,600
32 0.003 0.24 0.37 NN >10,000>\!10{,}000  |  Lang 7,400  |  Koop 5,600
 $_{\mbox{1}}A$R, criterion lDDT<0.65<0.65; Lang uses s=1s{=}1, $\Delta t{=}2{\times}10^{-9}$ (calibrated), $\tau{=}0$, clip c=10c{=}10.
We select $d_{z}{=}16$ in the main runs for a good trade-off between training cost and downstream accuracy.
A.7.2 (ii) Pooling width $H$ with W=2W{=}2 (so $d_{\text{flat}}{=}2H$)
Larger pooled latents reduce decoder MSE, but very large $d_{\text{flat}}$ slightly degrades flexibility/stability, likely by making the latent stiffer to traverse with simple propagators.
[…]
20  (40) 0.43 0.31 0.49 7,100 5,400
50  (100) 0.24 0.26 0.39 7,400 5,600
80  (160) 0.21 0.30 0.46 7,000 5,500
We therefore adopt H=50H{=}50 (i.e., $d_{\text{flat}}{=}100$) for all main results.
A.7.3 (iii) Denoiser/Langevin settings
We ablate the diffusion schedule for the denoiser (which controls score quality), the score query level $\tau$, the clipping threshold, and the SDE step size $\Delta t$ (with stride fixed to 1). Trends:
[…]
  • Clipping. Max-norm clip at c=10c{=}10 stabilized occasional spikes without biasing small updates; removing clip reduced the Langevin failure time from 7,4007{,}400 to 6,6006{,}600 frames; c=5c{=}5 was slightly over-damped (failure 7,2007{,}200).
[…]
Summary.
Bigger $d_{z}$ and pooled width reduce reconstruction losses as expected, but overly large pooled latents slightly worsen flexibility and stability. A well-trained denoiser with $\tau{=}0$, clip c=10c{=}10, and $\Delta t$ calibrated by per-frame RMSD matching yields the best side-chain fidelity and stable long rollouts on $_{\mbox{1}}A$R.
A.8 Supplementary Figures
Figure 5: Representative structural snapshots from latent rollouts. Top: Alanine–dipeptide conformers sampled from a Koopman rollout show backbone dihedral changes across frames. Bottom:  $_{\mbox{2}}A$Rsnapshots from an Langevin rollout; the dashed circle highlights the intracellular end of TM6, which moves outward over time—a hallmark of GPCR activation.
A.9 Supplementary Tables
Table 10: Summary of notation used in the description of the propagator models.
Symbol Description
$z_{t}\in\mathbb{R}^{d}$ Latent space vector at time $t$.
$\mathbf{X},\mathbf{Y}\in\mathbb{R}^{d\times(M-1)}$ Snapshot matrices with columns as time: $\mathbf{X}=[z_{0},\dots,z_{M-2}]$, $\mathbf{Y}=[z_{1},\dots,z_{M-1}]$.
$A\in\mathbb{R}^{d\times d}$ Koopman linear operator.
$f_{\theta}$ Autoregressive NN with parameters $\theta$.
$p(z)$ Equilibrium probability distribution of the latent variable $z$.
$s(z)=\nabla_{z}\log p(z)$ Score function of the equilibrium distribution.
$\epsilon_{\theta}(z_{t},\tau)$ Pre-trained LD-FPG diffusion model (denoiser).
$\sigma_{\tau}$ Noise schedule standard deviation from the diffusion model at step $\tau$.
$\Delta t$ Integration time step for Langevin dynamics.
$T$ Temperature parameter for Langevin dynamics.
$\eta_{t}$ Stochastic noise term, $\eta_{t}\sim\mathcal{N}(0,\sigma_{\eta}^{2}I)$ (scale $\sigma_{\eta}$ specified per experiment).
$d_{z}$ Per-atom encoder embedding width (ChebNet channels).
$d_{\text{flat}}$ Flattened pooled latent dimension after $H\times W$ pooling (here $50\times 2\Rightarrow d_{\text{flat}}=100$).
$d$ Shorthand for $d_{\text{flat}}$ (the flattened pooled latent dimension).
Table 11: Glossary of Biophysical Terms. Definitions of key concepts from molecular dynamics and structural biology used in this work, tailored for a machine learning audience.
Term Description for an ML Audience
Collective Variable (CV) A low-dimensional function of atomic coordinates (e.g., a distance or angle) designed to capture a specific, slow dynamic process like protein folding. Traditional simulation methods often require pre-defining good CVs; our work uses a learned latent space to discover them automatically.
Potential of Mean Force (PMF) Essentially, an effective "energy landscape" for a molecule in solution. Lower values correspond to more probable (stable) conformations. It is the target distribution that score-guided Langevin dynamics aims to sample from, often visualized as a "free-energy surface".
Dihedral Angles ($\phi,\psi,\chi$) Rotational angles around covalent bonds that define the geometry of a molecule. $\phi$ (phi), $\psi$ (psi): Define the rotation of the protein backbone. $\chi$ (chi): Define the rotation of the amino acid side-chains. Their statistical distributions are a sensitive measure of structural fidelity.
Ramachandran Plot A 2D plot of the backbone dihedral angles ($\phi,\psi$). Certain regions of this plot are "allowed" based on steric constraints, leading to characteristic high-probability basins that correspond to stable secondary structures like alpha-helices and beta-sheets.
Rotamer A discrete, low-energy, and therefore highly probable conformation of a protein’s side-chain, defined by its set of $\chi$ angles. A key test for generative models is whether they can reproduce the correct statistical distribution of these rotameric states.
GPCR G protein-coupled receptor. A large and important family of transmembrane proteins that act as cellular signal transducers. They are highly dynamic and switch between different functional states (e.g., inactive, active), making them an ideal and challenging test system for dynamic models.
lDDT local Distance Difference Test. A metric for assessing the quality of a protein structure prediction by evaluating how well local inter-atomic distances are preserved relative to a reference structure. Unlike RMSD, it is less sensitive to global rotations and more focused on local geometric accuracy.
RMSF Root-Mean-Square Fluctuation. For each atom, this metric calculates the standard deviation of its position over time in a simulation trajectory. It measures the "amplitude of motion" or flexibility of different parts of the protein.